OpenStudy (anonymous):

(3,-3),y=x+5 Write an equation of the line that passes through the given point and is perpendicular to the given line.

OpenStudy (anonymous):

x=y

OpenStudy (anonymous):

dont you have to make the given slope opposite

OpenStudy (anonymous):

sorry x=-y

OpenStudy (anonymous):

sritama idont think thats right

OpenStudy (anonymous):

why not -x

OpenStudy (anonymous):

perpendicular means if the first line's gradient is n, the other line's gradient is -1/n so the line you are looking for has the gradient of -1/1 which is just -1 (y-(-3))/(x-3)=-1 y+3=-x+3 y=-x+6

OpenStudy (anonymous):

oops y+3=-x+3 y=-x silly mistake, sorry

OpenStudy (anonymous):

hmmmm

OpenStudy (anonymous):

 Draw the picture. It all becomes clear.